Use four iterations of Romberg integration to estimate π = Jo 22 +1
Answer #1

=fade -0, b= 1, flm- Rus = [f@14f0J = 46) - (4 + 2) = 3. R21 - 4 (40)+24(4) +-600] IT -à (4+2.16 +2) = 1+ 2+ Ź = 1+ 8 +3 = 10ARALLARHOLM Rai to (50) + 2(363) +() + f(a) ** (2) *fl@ +4(9)+f@)) + fu) (+ 2 (25 4 9 2 27 101869913599 32449916200 R2,2 = R2R3,3 = R3, 2 + 1 R3,2 R2,2) 6677 2125 Ra3 = R4,2 + 1 (R 2 - R3,2) 127430581361 40662396250 RGA = R4,3 + 3 (R4,3 R3.3) 4014062